# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

### v1

@@ 1. $\mathcal{H}\leftarrow\emptyset$
2. $seed\leftarrow\emptyset$ $minHS(\mathcal{H})$
3. ⊧̸$\emptyset\not\models a∅ ⊧$ $SAT(\epsilon\cup\{\lnot a\})$
4. $\mathcal{C}\leftarrow\{C_{1}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
@@ 5. $\mathcal{H}\leftarrow\{\{C_{1}\}\}$
6. $seed\leftarrow\{C_{1}\}$ $minHS(\mathcal{H})$
7. $\{a\vee b\}\not\models a$ $SAT(\epsilon\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{2}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
@@ 9. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\}\}$
10. $seed\leftarrow\{C_{1},C_{2}\}$ $minHS(\mathcal{H})$
11. $\{a\vee b,\lnot b\vee c\}\not\models a$ $SAT(\epsilon\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{3}\}$ MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed $seed$
@@ 13. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\},\{C_{3}\}\}$
14. $seed\leftarrow\{C_{1},C_{2},C_{3}\}$ $minHS(\mathcal{H})$
15. $\{a\vee b,\lnot b\vee c,\lnot c\}\models a$$\lnot SAT(\epsilon\cup\{\lnot a\})$
16. $Return$$\{a\vee b,\lnot b\vee c,\lnot c\}$ minimum size monolithic explanation for $a$ from $\mathrm{KB}$
Algorithm 1 describes the main steps of our approach. $\mathcal{H}$ is a collection of sets, where each set corresponds to an MCS on $\mathrm{KB}$. At the beginning, it is initialized with the empty set (line 1). Each MCS in $\mathcal{H}$ is represented as the set of the indexes of the formulae in it. $\mathcal{H}$ stores the MCSes computed so far. At each step, a minimal hitting set on $\mathcal{H}$ is computed (line 1). In line 1, the formulae induced by the computed minimal hitting set is stored in $\epsilon$. Then, $\epsilon\cup\{\lnot\varphi\}$ is evaluated for satisfiability (line 1). If $\epsilon\cup\{\lnot\varphi\}$ is inconsistent, then $\epsilon$ is a monolithic explanation of minimum size and the algorithm returns $\epsilon$. If instead $\epsilon\cup\{\lnot\varphi\}$ is consistent, then it means that $\epsilon\not\models\varphi$ and the algorithm continues in line 1. The computation of an MCS of this kind can be performed via standard MCS procedures (?), using the set of formulae indexed by the seed as the starting formula to extend. Since $\varphi$ is set to hard (line 1), the returned MCS $\mathcal{C}$ is guaranteed to be contained in $\mathrm{KB}$. Due to the hitting set duality relation, we will also have $\epsilon\subseteq\mathrm{KB}$. Finally, notice that the procedure $getMCS$ always reports a new MCS because, by construction, we have $seed\subseteq\mathrm{KB}\setminus\mathcal{C}$. In fact, the $seed$ contains at least one formula for each previously computed MCS and, thus, $seed\cap\mathcal{C}=\emptyset$ (i.e., at least one formula for each previously computed MCS is not in $\mathcal{C}$).
@@ 12                   $\epsilon^{-}\leftarrow\texttt{getMCS}((\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s})$
17             $\mathcal{C}\leftarrow\texttt{getMCS}(seed,\mathrm{KB}_{a}^{h}\cup\{\lnot\varphi^{h}\}\cup\mathrm{KB}_{\alpha}^{s})$
@@ 18             $\mathbb{R}\leftarrow\mathbb{R}\cup\{\mathcal{C}\}$
Algorithm 2 model-reconciling-explanation$(\mathrm{KB}_{\alpha},\mathrm{KB}_{a}^{h},\varphi)$
@@ 2. $\mathrm{KB}_{\alpha}^{h}\leftarrow\mathrm{KB}_{\alpha}\cap\mathrm{KB}_{h}=\{C_{3}\}$
4. $E^{-}\leftarrow\{D_{1}\}$ MCS computed on $(\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha})^{s}\cup\mathrm{KB}_{\alpha}^{h}$
@@ 5. $\mathrm{KB}_{h}\leftarrow\{D_{1},D_{2}\}\setminus\{D_{1}\}=\{D_{2}\}$
6. $seed\leftarrow\emptyset$ $minHS(\mathcal{R})$
7. $\{\lnot c\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{1}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 9. $\mathcal{R}\leftarrow\{\{C_{1}\}\}$
10. $seed\leftarrow\{C_{1}\}$ $minHS(\mathcal{R})$
11. $\{\lnot c,a\vee b\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{2},C_{4}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 13. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\}\}$
14. $seed\leftarrow\{C_{1},C_{4}\}$ $minHS(\mathcal{R})$
15. $\{\lnot c,a\vee b,\lnot b\vee d\}\not\models a$ $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
16. $\mathcal{C}\leftarrow\{C_{2},C_{5}\}$ MCS computed on $\mathrm{KB}_{a}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 17. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\},\{C_{2},C_{5}\}\}$
18. $seed\leftarrow\{C_{1},C_{2}\}$ $minHS(\mathcal{R})$
19. $\{\lnot c,a\vee b,\lnot b\vee c\}\models a$$\lnot SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
20. $\epsilon^{-}\leftarrow\{D_{1}\}$ MCS computed on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$
21. $Return$$\langle\{C_{1},C_{2}\},\{D_{1}\}\rangle$ model reconciling explanation for $a$ from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$
@@ 4.1.1 Model Reconciling Explanations
Algorithm 2 describes the pseudocode of our approach. At the beginning of the algorithm, we initialize $\mathbb{R}$ to the null set (line 2). $\mathbb{R}$ is used to store the MCSes, which acts as a mediator between $\mathrm{KB}_{\alpha}$ and $\mathrm{KB}_{h}$. Lines 2-2 are used to specify which clauses of $\mathrm{KB}_{\alpha}$ will be treated as hard and soft constraints, respectively. We then check if $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent (line 2). This is important in order to avoid the possibility of finding subsets $\epsilon^{+}$ that explain why $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent instead of the target explanandum. In case $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent, we preprocess $\mathrm{KB}_{h}$ by removing from $\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha}$ a minimal set of formulae causing the conflict (i.e., an MCS) (lines 2-2), where $E^{-}$ stores the set of potential formulae $\epsilon^{-}$ to retract. The reconciliation procedure starts in line 2. The algorithm proceeds iteratively by computing a minimal hitting set on $\mathbb{R}$ and then testing for satisfiability the formulae $\epsilon^{+}$ (lines 2-2). The test checks whether adding $\epsilon^{+}$ to $\mathrm{KB}_{h}$ is sufficient for entailing $\varphi$. If $\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot\varphi\}$ is unsatisfiable, then $\mathrm{KB}_{h}\cup\epsilon^{+}\models\varphi$. In that case, the algorithm then checks whether $\mathrm{KB}_{h}\cup\epsilon^{+}\cup E^{-}$ is inconsistent, and if it is, it computes an MCS $e^{-}$ on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$ (lines 2-2). The model reconciling explanation $\langle\epsilon^{+},\epsilon^{-}\rangle$ is then returned in line 2. Otherwise, the algorithm continues in line 2, where a new MCS is computed and added to $\mathbb{R}$ [^12] Note that the algorithm is complete as it is based on Algorithm 1, which is complete.
@@ 4.2.1 Monolithic Explanations
Consider an explanandum $\varphi$ and a belief base $\mathcal{B}$. First, notice that if we assume that the classical projection of $\mathcal{B}$ entails the explanandum $\varphi$, that is $\mathcal{B}^{\downarrow w}\models\varphi$, then Algorithm 1 can directly be applied on $\mathcal{B}^{\downarrow w}$ and $\varphi$ [^13] Recall that the classical projection of belief base $\mathcal{B}$ is the unweighted version of the set of formulae from $\mathcal{B}$. In that case, Algorithm 1 guarantees to find a monolithic explanation with maximum explanatory gain, since we know from Proposition 3 that explanatory gain achieves its maximum value for $\varphi$ when the monolithic explanation entails $\varphi$. Nevertheless, this does not guarantee that the monolithic explanation will be the most-preferred one, that is, the one with the highest explanatory power (Definition 13).
Obviously, a straightforward way of computing a most-preferred monolithic explanation is to use Algorithm 1 to enumerate all possible monolithic explanations for $\varphi$, and return the one that has the highest probability, which corresponds to the one with the highest explanatory power. But enumerating through all possible monolithic explanations and computing their probabilities can be computationally prohibited, as even extracting a smallest monolithic explanation is in $FP^{\Sigma^{p}_{2}}$ (?) and computing the probability of a formula is #$\#P$-complete (?, ?). We can, however, account for this high computational complexity by seeking for a monolithic explanation that is guaranteed to have a probability above a certain threshold.
@@
[…]
@@
Consider the belief base =$\mathcal{B}=\{(a,1),(\lnot a\vee b,3),(c,2),(\lnot c\vee b,1)\}$ and explanandum $b$. The two monolithic explanations for $b$ from $\mathcal{B}$ that entail $b$ are $\tilde{\epsilon}_{1}=\{a,\lnot a\vee b\}$ and $\tilde{\epsilon}_{2}=\{c,\lnot c\vee b\}$, where $P(\tilde{\epsilon}_{1})=0.64$ and $P(\tilde{\epsilon}_{2})=0.77$. Notice that there are four possible worlds in which $b$ is true: $\omega_{1}=\{a=T,b=T,c=T\},$ $\omega_{2}=\{a=F,b=T,c=T\}$, $\omega_{3}=\{a=T,b=T,c=F\}$, and $\omega_{4}=\{a=F,b=T,c=F\}$, where $P(\omega_{1})=0.57$, $P(\omega_{2})=0.20$, $P(\omega_{3})=0.07$, and $P(\omega_{4})=0.02$. The maximum number of intersections that entail $b$ is $k=2$ (i.e., $I_{2}=\{\lnot a\vee b,c,\lnot c\vee b\}$). Indeed, $\tilde{\epsilon}_{2}\subseteq I_{2}$ and $P(\tilde{\epsilon}_{2})=0.77=P(\omega_{1})+P(\omega_{2})$. Finally, notice how $\tilde{\epsilon}_{2}$ is also the most-preferred monolithic explanation for $b$ from $\mathcal{B}$; for $\gamma=0.5$, $\mathrm{EP}(\tilde{\epsilon}_{2},b)=0.57>\mathrm{EP}(\tilde{\epsilon}_{1},b)=0.50$.
@@
The proof follows directly from Lemma 2.
@@
The utility of a $k$-bounded monolithic explanation in computing probabilistic monolithic explanations can be described as follows. If we take the top $k$ most-probable worlds in which the explanandum $\varphi$ is true, then we can prune the search space of possible monolithic explanations by taking the intersection of those worlds and checking if it entails $\varphi$ – if it does, then we know that at least one monolithic explanation must be true in that world with probability at least equal to the sum of the probabilities of these top $k$ worlds. Building on this, we now present an algorithm for computing $k$-bounded monolithic explanations for $\varphi$ from $\mathcal{B}$, where we use Algorithm 1 as our core monolithic explanation generation engine.

### v2

@@ 1. $\mathcal{H}\leftarrow\emptyset$
2. $seed\leftarrow\emptyset$ # $minHS(\mathcal{H})$
3. $\emptyset\not\models a$ # $SAT(\epsilon\cup\{\lnot a\})$
4. $\mathcal{C}\leftarrow\{C_{1}\}$ # MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed seedseed
@@ 5. $\mathcal{H}\leftarrow\{\{C_{1}\}\}$
6. $seed\leftarrow\{C_{1}\}$ # $minHS(\mathcal{H})$
7. $\{a\vee b\}\not\models a$ # $SAT(\epsilon\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{2}\}$ # MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed seedseed
@@ 9. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\}\}$
10. $seed\leftarrow\{C_{1},C_{2}\}$ # $minHS(\mathcal{H})$
11. $\{a\vee b,\lnot b\vee c\}\not\models a$ # $SAT(\epsilon\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{3}\}$ # MCS computed on $KB^{s}\cup\{\lnot a^{h}\}$ starting with the seed seedseed
@@ 13. $\mathcal{H}\leftarrow\{\{C_{1}\},\{C_{2}\},\{C_{3}\}\}$
14. $seed\leftarrow\{C_{1},C_{2},C_{3}\}$ # $minHS(\mathcal{H})$
15. $\{a\vee b,\lnot b\vee c,\lnot c\}\models a$ # $\lnot SAT(\epsilon\cup\{\lnot a\})$
16. ReturnReturn $\{a\vee b,\lnot b\vee c,\lnot c\}$ # minimum size monolithic explanation for $a$ from $\mathrm{KB}$
Algorithm 1 describes the main steps of our approach. $\mathcal{H}$ is a collection of sets, where each set corresponds to an MCS on $\mathrm{KB}$. At the beginning, it is initialized with the empty set (line 1). Each MCS in $\mathcal{H}$ is represented as the set of the indexes of the formulae in it. $\mathcal{H}$ stores the MCSes computed so far. At each step, a minimal hitting set on $\mathcal{H}$ is computed (line 1). In line 1, the formulae induced by the computed minimal hitting set is stored in $\epsilon$. Then, $\epsilon\cup\{\lnot\varphi\}$ is evaluated for satisfiability (line 1). If $\epsilon\cup\{\lnot\varphi\}$ is inconsistent, then $\epsilon$ is a monolithic explanation of minimum size and the algorithm returns $\epsilon$. If instead $\epsilon\cup\{\lnot\varphi\}$ is consistent, then it means that $\epsilon\not\models\varphi$ and the algorithm continues in line 1. The computation of an MCS of this kind can be performed via standard MCS procedures (?), using the set of formulae indexed by the seed as the starting formula to extend. Since $\varphi$ is set to hard (line 1), the returned MCS $\mathcal{C}$ is guaranteed to be contained in $\mathrm{KB}$. Due to the hitting set duality relation, we will also have $\epsilon\subseteq\mathrm{KB}$. Finally, notice that the procedure $getMCS$ always reports a new MCS because, by construction, we have $seed\subseteq\mathrm{KB}\setminus\mathcal{C}$. In fact, the seedseed contains at least one formula for each previously computed MCS and, thus, $seed\cap\mathcal{C}=\emptyset$ (i.e., at least one formula for each previously computed MCS is not in $\mathcal{C}$).
@@ 12       $\epsilon^{-}\leftarrow\texttt{getMCS}((\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s})$
17    $\mathcal{C}\leftarrow\texttt{getMCS}(seed,\mathrm{KB}_{\alpha}^{h}\cup\{\lnot\varphi^{h}\}\cup\mathrm{KB}_{\alpha}^{s})$
@@ 18    $\mathbb{R}\leftarrow\mathbb{R}\cup\{\mathcal{C}\}$
[…]
@@ 2. $\mathrm{KB}_{\alpha}^{h}\leftarrow\mathrm{KB}_{\alpha}\cap\mathrm{KB}_{h}=\{C_{3}\}$
4. $E^{-}\leftarrow\{D_{1}\}$ # MCS computed on $(\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha})^{s}\cup\mathrm{KB}_{\alpha}^{h}$
@@ 5. $\mathrm{KB}_{h}\leftarrow\{D_{1},D_{2}\}\setminus\{D_{1}\}=\{D_{2}\}$
6. $seed\leftarrow\emptyset$ # $minHS(\mathcal{R})$
7. $\{\lnot c\}\not\models a$ # $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
8. $\mathcal{C}\leftarrow\{C_{1}\}$ # MCS computed on $\mathrm{KB}_{\alpha}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 9. $\mathcal{R}\leftarrow\{\{C_{1}\}\}$
10. $seed\leftarrow\{C_{1}\}$ # $minHS(\mathcal{R})$
11. $\{\lnot c,a\vee b\}\not\models a$ # $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
12. $\mathcal{C}\leftarrow\{C_{2},C_{4}\}$ # MCS computed on $\mathrm{KB}_{\alpha}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 13. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\}\}$
14. $seed\leftarrow\{C_{1},C_{4}\}$ # $minHS(\mathcal{R})$
15. $\{\lnot c,a\vee b,\lnot b\vee d\}\not\models a$ # $SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
16. $\mathcal{C}\leftarrow\{C_{2},C_{5}\}$ # MCS computed on $\mathrm{KB}_{\alpha}^{h}\cup\mathrm{KB}_{\alpha}^{s}\cup\{\lnot a^{h}\}$
@@ 17. $\mathcal{R}\leftarrow\{\{C_{1}\},\{C_{2},C_{4}\},\{C_{2},C_{5}\}\}$
18. $seed\leftarrow\{C_{1},C_{2}\}$ # $minHS(\mathcal{R})$
19. $\{\lnot c,a\vee b,\lnot b\vee c\}\models a$ # $\lnot SAT(\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot a\})$
20. $\epsilon^{-}\leftarrow\{D_{1}\}$ # MCS computed on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$
21. ReturnReturn $\langle\{C_{1},C_{2}\},\{D_{1}\}\rangle$ # model reconciling explanation for $a$ from $\mathrm{KB}_{\alpha}$ for $\mathrm{KB}_{h}$
@@ 6.1.1 Model Reconciling Explanations
Algorithm 2 describes the pseudocode of our approach. At the beginning of the algorithm, we initialize $\mathbb{R}$ to the null set (line 2). $\mathbb{R}$ is used to store the MCSes, which acts as a mediator between $\mathrm{KB}_{\alpha}$ and $\mathrm{KB}_{h}$. Lines 2-2 are used to specify which clauses of $\mathrm{KB}_{\alpha}$ will be treated as hard and soft constraints, respectively. We then check if $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent (line 2). This is important in order to avoid the possibility of finding subsets $\epsilon^{+}$ that explain why $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent instead of the target explanandum. In case $\mathrm{KB}_{h}\cup\mathrm{KB}_{\alpha}$ is inconsistent, we preprocess $\mathrm{KB}_{h}$ by removing from $\mathrm{KB}_{h}\setminus\mathrm{KB}_{\alpha}$ a minimal set of formulae causing the conflict (i.e., an MCS) (lines 2-2), where $E^{-}$ stores the set of potential formulae $\epsilon^{-}$ to retract. The reconciliation procedure starts in line 2. The algorithm proceeds iteratively by computing a minimal hitting set on $\mathbb{R}$ and then testing for satisfiability the formulae $\epsilon^{+}$ (lines 2-2). The test checks whether adding $\epsilon^{+}$ to $\mathrm{KB}_{h}$ is sufficient for entailing $\varphi$. If $\mathrm{KB}_{h}\cup\epsilon^{+}\cup\{\lnot\varphi\}$ is unsatisfiable, then $\mathrm{KB}_{h}\cup\epsilon^{+}\models\varphi$. In that case, the algorithm then checks whether $\mathrm{KB}_{h}\cup\epsilon^{+}\cup E^{-}$ is inconsistent, and if it is, it computes an MCS $e^{-}$ on $(\mathrm{KB}_{h}\cup\epsilon^{+})^{h}\cup(E^{-})^{s}$ (lines 2-2). The model reconciling explanation $\langle\epsilon^{+},\epsilon^{-}\rangle$ is then returned in line 2. Otherwise, the algorithm continues in line 2, where a new MCS is computed and added to $\mathbb{R}$. [^15] Note that the algorithm is complete as it is based on Algorithm 1, which is complete.
@@ 6.2.1 Monolithic Explanations
Consider an explanandum $\varphi$ and a belief base $\mathcal{B}$. First, notice that if we assume that the classical projection of $\mathcal{B}$ entails the explanandum $\varphi$, that is $\mathcal{B}^{\downarrow w}\models\varphi$, then Algorithm 1 can directly be applied on $\mathcal{B}^{\downarrow w}$ and $\varphi$. [^16] Recall that the classical projection of belief base $\mathcal{B}$ is the unweighted version of the set of formulae from $\mathcal{B}$. In that case, Algorithm 1 guarantees to find a monolithic explanation with maximum explanatory gain, since we know from Proposition 3 that explanatory gain achieves its maximum value for $\varphi$ when the monolithic explanation entails $\varphi$. Nevertheless, this does not guarantee that the monolithic explanation will be the most-preferred one, that is, the one with the highest explanatory power (Definition 14).
Obviously, a straightforward way of computing a most-preferred monolithic explanation is to use Algorithm 1 to enumerate all possible monolithic explanations for $\varphi$, and return the one that has the highest probability, which corresponds to the one with the highest explanatory power. But enumerating through all possible monolithic explanations and computing their probabilities can be computationally prohibitive, as even extracting a smallest monolithic explanation is in $FP^{\Sigma^{p}_{2}}$ (?) and computing the probability of a formula is #P\#P-complete (?, ?). We can, however, account for this high computational complexity by seeking for a monolithic explanation that is guaranteed to have a probability above a certain threshold.
@@
Let $\tilde{E}(\varphi)$ be the set of all monolithic explanations for explanandum $\varphi$ from belief base $\mathcal{B}$. Let $\Omega(\varphi)=\{\omega_{1},\ldots,\omega_{n}\}$ be the set of possible worlds in which $\varphi$ is true, where $P(\omega_{1})\geq P(\omega_{2})\geq\ldots\geq P(\omega_{n})$. Also let $I_{k}=\overset{k}{\underset{i=1}{{{\bigcap}}}}\{\phi\>|\>\phi\in\mathcal{B}^{\downarrow w},\omega_{i}\models\phi\}$ be the intersection of formulae that are true in worlds $\omega_{1}$ to $\omega_{k}$. We say that $\tilde{\epsilon}\in\tilde{E}(\varphi)$ is a $k$-bounded monolithic explanation for $\varphi$ from $\mathcal{B}$, with lower bound $P(\tilde{\epsilon})\geq\overset{k}{\underset{i=1}{{{\sum}}}}P(\omega_{i})$, if and only if $\tilde{\epsilon}\subseteq I_{k}$.
@@
Consider the belief base $\mathcal{B}=\{(a,1),(\lnot a\vee b,3),(c,2),(\lnot c\vee b,1)\}$ and explanandum $b$. The two monolithic explanations for $b$ from $\mathcal{B}$ that entail $b$ are $\tilde{\epsilon}_{1}=\{a,\lnot a\vee b\}$ and $\tilde{\epsilon}_{2}=\{c,\lnot c\vee b\}$, where $P(\tilde{\epsilon}_{1})=0.64$ and $P(\tilde{\epsilon}_{2})=0.77$. Notice that there are four possible worlds in which $b$ is true: $\omega_{1}=\{a=T,b=T,c=T\},$ $\omega_{2}=\{a=F,b=T,c=T\}$, $\omega_{3}=\{a=T,b=T,c=F\}$, and $\omega_{4}=\{a=F,b=T,c=F\}$, where $P(\omega_{1})=0.57$, $P(\omega_{2})=0.20$, $P(\omega_{3})=0.07$, and $P(\omega_{4})=0.02$. The maximum number of intersections that entail $b$ is $k=2$ (i.e., $I_{2}=\{\lnot a\vee b,c,\lnot c\vee b\}$). Indeed, $\tilde{\epsilon}_{2}\subseteq I_{2}$ and $P(\tilde{\epsilon}_{2})=0.77=P(\omega_{1})+P(\omega_{2})$. Finally, notice how $\tilde{\epsilon}_{2}$ is also the most-preferred monolithic explanation for $b$ from $\mathcal{B}$; for $\gamma=0.5$, $\mathrm{EP}(\tilde{\epsilon}_{2},b)=0.57>\mathrm{EP}(\tilde{\epsilon}_{1},b)=0.50$.
@@
The proof follows directly from Lemma 2. ∎
@@
The utility of a $k$-bounded monolithic explanation in computing probabilistic monolithic explanations can be described as follows. If we take the top $k$ most-probable worlds in which the explanandum $\varphi$ is true, then we can prune the search space of possible monolithic explanations by taking the intersection of those worlds and checking if it entails $\varphi$—if it does, then we know that at least one monolithic explanation must be true in that world with probability at least equal to the sum of the probabilities of these top $k$ worlds. Building on this, we now present an algorithm for computing $k$-bounded monolithic explanations for $\varphi$ from $\mathcal{B}$, where we use Algorithm 1 as our core monolithic explanation generation engine.